# Fleurs-SLU: A Massively Multilingual Benchmark for Spoken Language Understanding

### v2

@@ 6 Conclusion
We introduce Fleurs-SLU, a multilingual SLU benchmark for semantic speech classification across 102 languages and multiple-choice question answering from spoken paragraphs in 92 languages. Using Fleurs-SLU, we evaluate massively multilingual speech models in both end-to-end speech classification and a cascaded approach that combines initial speech-to-text transcription and subsequent text-based classification with LLMs. Our findings indicate that, while cascaded systems remain the most robust option, multilingual speech encoders can achieve competitive performance when adequately pre-trained. Furthermore, we observe a strong correlation between strong multilingual SLU and both the robustness of multilingual ASR and the effectiveness of cross-modal speech translation to English text. This suggests that multilingual SLU and multilingual ASR can be mutually beneficial. We hope that our findings inspire future work towards developing more efficient multilingual speech encoders that are jointly pre-trained for both multilingual ASR and SLU to close the performance gap between end-to-end speech classification and cascaded approaches.
Limitations
Our work has two primary limitations. First, the selection of models in CS can be further expanded. In particular, additional evaluations on CS that combine transcription models with prompting 70B LLMs would provide valuable insights given the growing interest in instruction fine-tuned LLMs. Moreover, assessing speech-LLMs on both tasks would further illuminate multimodal alignment in massively multilingual scenarios. Notably, LLM2Vec on Llama 3.1 8B, which underwent extensive task fine-tuning on 63K instances, represents an exceptionally strong language model backbone for CS. Compared to concurrent work, our models would have even outperformed prompting 70B LLMs on speech transcriptions in Belebele Costa-jussà et al. (2024). Additionally, current speech LLMs rely on Whisper, which is inherently limited to 30-second speech segments. This limitation hinders their application in Belebele-Fleurs, where the average speech length is 47 seconds (cf. Table 2). Our compute infrastructure was unable to support prompting such large LLMs on long contexts, exceeding half a minute of speech combined with corresponding questions and answers. We therefore leave these evaluations for future work.
Furthermore, although our findings indicate a clear link between robust multilingual SLU and strong multilingual ASR, future research must more carefully isolate the mechanisms underlying this interaction. Such investigations should clarify the conditions (e.g., training objectives, data and language mixtures) under which SLU most effectively contributes to more robust multilingual ASR. Nonetheless, our analysis strongly suggests that multilingual SLU and multilingual ASR are mutually beneficial.
Acknowledgments
We use AI assistance, specifically chatGPT o3-mini, to refine both the writing and the tables in our manuscript.
@@ A.1 Further Experimental Details
Code. The code to reproduce the Fleurs-SLU benchmark is available at: https://github.com/fdschmidt93/fleurs-slu.

### v3

@@ 6 Conclusion
We introduce Fleurs-SLU, a multilingual SLU benchmark for semantic speech classification across 102 languages and multiple-choice question answering from spoken paragraphs in 92 languages. Using Fleurs-SLU, we evaluate massively multilingual speech models in both end-to-end speech classification and a cascaded approach that combines initial speech-to-text transcription and subsequent text-based classification with LLMs. Our findings indicate that, while cascaded systems remain the most robust option, multilingual speech encoders can achieve competitive performance when adequately pre-trained. Moreover, speech-LLMs can yield state-of-the-art performance on par with the best targeted SLU pipelines when appropriately aligned for multilingual spoken language instruction following. Moreover, speech-LLMs can achieve state-of-the-art performance on par or better than the best targeted SLU pipelines when properly aligned for multilingual spoken language instruction following. Furthermore, we observe a strong correlation between strong multilingual SLU and both the robustness of multilingual ASR and the effectiveness of cross-modal speech translation to English text. This suggests that multilingual SLU and multilingual ASR can be mutually beneficial. We hope that our findings inspire future work towards developing more efficient multilingual speech encoders that are jointly pre-trained for both multilingual ASR and SLU to close the performance gap between end-to-end speech classification and cascaded approaches.
@@ A.1 Further Experimental Details
Prompting Qwen 2.5 7B-Omni & Gemini 2-Flash.
For the multimodal speech-LLMs Qwen 2.5 7B-Omni and Gemini 2-Flash, we apply the following utterance pre-processing. Each audio segment is normalized to a target root mean square (RMS) level of 0.07 to ensure consistent loudness across samples and reduce variability due to recording conditions. Specifically, we first normalize the loudness of individual utterances; for Belebele-Fleurs, we additionally normalize the loudness after concatenation. The models are then prompted using the task- and modality-specific templates detailed below.